The quadratic function which takes the value 41 when x=-2 and the value 20 at x=5 is y=ax^2-bx+c.Need help in finding values for a,b,c
Answer by Boreal(15235)

I would use ax^2+bx+c as the function, which will allow the b to be both positive or negative.
4a-2b+c=41
25a+5b+c=20
subtract the second from the first
-21a-7b=21
divide by 7, both sides
-3a-b=3
b=-3a-3=-3(a+1)
go back to original equations
4a+6(a+1)+c=41
25a-15(a+1)+c=20
subtract again
-21a+9(a+1)=21
-21a+9a+9=21
-12a=12
a=-1
b=0
c=45
y=-x^2+45